the circle x^2+y^2+4x-7y+12=0 cuts an intercept on y axis equal to ???

Grade:11

2 Answers

Shubham kumawat
6 years ago
Length of y-intercept =2√(f2 -c) f=7/2, c=12 2√(49/4 – 12) =1 Or we can also find by finding radius and the distance of centre from y axis and then apply pythogorus .
Arun
6 years ago
Dear Prasad
 
Y intercept = 2 sqrt (f²-c) 
Here f = -7/2 , c = 12
 
Now I am sure you can do it.
 
Regards
Arun (askIITians forum expert)